OpenStudy (idealist10):

Find the general solution of xy"+(2-2x)y'+(x-2)y=0, given that y1=e^x satisfies the complementary equation.

OpenStudy (idealist10):

y=ue^x y'=u'e^(x)+ue^(x) y"=u"e^(x)+2u'e^(x)+ue^(x) Subbing: xy"+(2-2x)y'+(x-2)y=u"xe^(x)+2u'e^(x)=0 \[u''+\frac{ 2 }{ x }u'=0\] z=u' Subbing: \[z'+\frac{ 2 }{ x }z=0\] the integrating factor is x^2 \[x^2(z'+\frac{ 2 }{ x }z)=0\] \[\frac{ d }{ dx }(x^2z)=0\] \[x^2z=C\] \[z=u'=\frac{ C }{ x^2 }\]

OpenStudy (idealist10):

integrate, \[u=-\frac{ C _{1} }{ x }+C _{2}\] So \[y=e ^{x}(-\frac{ C _{1} }{ x }+C _{2})\]

OpenStudy (idealist10):

Is my answer right? Can anyone confirm?

OpenStudy (idealist10):

@Kainui @Luigi0210 @Preetha

OpenStudy (kainui):

What do you get when you plug it into the differential equation?